OpenStudy (daisyblake11):

please can someone solve: tan(x+pi)+2sin(x+pi)=0

Directrix (directrix):

I think you'll need these two formulas. The sine of the sum of two angles, and the tangent of the sum of two angles.

OpenStudy (mathmale):

Please look for ways in which you could simplify the given equation using the sum formulas mentioned above. Please show your work if you want feedback on it.

OpenStudy (daisyblake11):

i will try using the formulas

OpenStudy (daisyblake11):

still din't get it

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

tangent is periodic, with period \(\pi\) i.e \[\tan(x+\pi)=\tan(x)\] that might help

OpenStudy (misty1212):

also it might be useful to know that \[\sin(x+\pi)=-\sin(x)\]

OpenStudy (daisyblake11):

i dont get how tan(x+pi) = tanx . can you explain better

OpenStudy (misty1212):

i don't really know any other way to say it tangent is a periodic function, and the period is \(\pi\) that basically means that \[\tan(x+\pi)=\tan(x)\] just like sine and cosine have period \(2\pi\) so \[\sin(x+2\pi)=\sin(x)\]

OpenStudy (daisyblake11):

thank you

OpenStudy (misty1212):

how you are going to solve \[\tan(x)-2\sin(x)=0\] is anyones guess however

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (misty1212):

maybe you can do it "by inspection" 0 is a pretty obvious solution but there are more